# arXiv:2603.01180v1[astro-ph.HE] 01 Mar 2026

### 01 Mar 2026 v1

@@ Abstract
We study outflow from an unmagnetized, shocked accretion disk around a non-rotating super-massive black hole using multidimensional hydrodynamics simulation with radiative cooling. We aim to investigate whether such shocked accretion flow can launch sustained collimated bipolar outflow reaching out to thousands of gravitational radii even in the absence of magnetic field and if yes, what terminal velocity can they achieve. We present the results of a few simulations of geometrically thick accretion flow with increasing specific angular momentum on a vertically elongated cylindrical domain. We show that bipolar outflow from a region very close to the black hole is originating and propagating vertically out to our simulation domain boundary at around $\color[rgb]{0,0,0}2651$} Schwarzschild radius. The outflow attains a terminal velocity with a maximum value found to be $0.14c$ and the outflow rate depends on the angular momentum value of the accreting material. We also compute the self-Comptonized bremsstrahlung spectra for all the disk-jet runs.
@@ 1 Introduction
Many high-energy sources containing black hole show wind and/or jet (Mirabel and Rodríguez 1994; Stirling et al. 2001; Gallo 2010; Blandford et al. 2019). Whether the black hole is part of a binary system or located at the core of a galaxy, the qualitative structure of outflow is very much similar when it is present. The outflow originates from very close to the central object and extends from hundreds to thousands of gravitational radii. It is well understood that, unlike a normal star, a black hole itself does not eject matter. Rather, the gravitationally captured matter, forming an accretion disk around the black hole, produces the outflow. Given that the gravitational attraction is very strong close to a black hole, it is thus very surprising how a fraction of the infalling matter gets rid of this strong attraction and becomes unbounded.
@@
Outflow formation from the magnetized plasma of an accretion disk is mainly attributed to two mechanisms: Blandford-Znayek (Blandford and Znajek 1977) and Blandford-Pyne (Blandford and Payne 1982). It is of course not very surprising that the accreting material is magnetized. Magnetic field is very pervasive and wander along with plasma. In a binary system, matter can drag in magnetic field from the companion star. In the case of accretion onto a galaxy center, the infalling matter can drag magnetic field from nearby stars or from the interstellar medium. As this magnetized matter moves closer to the black hole, the magnetic field structure as well as the magnitude may change due to various physical processes such as geometric compression, dynamo effect, magnetic reconnection etc. This magnetic field can then help in shaping the accretion disk structure and launching outflow from the accretion disk in the manner described in these reference.
@@
However, magnetic field may not be the only driving force behind launching outflow. In principle, if a sufficiently strong force develops in a way that cancels the gravitational pull, matter can be driven away from the accretion disk. Thus, if the outward centrifugal force due to rotating matter becomes sufficiently strong, the effective potential barrier (Misner et al. 1973; Shapiro and Teukolsky 1983) may be so high that the matter may not be able to accrete onto the central black hole and be driven away. Also, a sufficiently strong pressure gradient force acting opposite to the gravitational attraction can drive material away. The pressure can be due to radiation or thermal. If the disk material and the radiation are interacting effectively, radiation pressure can act in that case (Icke 1989; Fukue 1996; Chattopadhyay et al. 2004; Vyas et al. 2015). On the other hand, if the cooling time scale is lower than the infall time scale, matter can become sufficiently hot and required thermal pressure gradient force can develop inside the accretion flow (e.g., Blandford and Begelman 1999; Soker 2023). If an accretion disk, having satisfied such conditions, additionally carries magnetic field, all the effects may add up to make the outflow launching favorable (e.g., Fukue 2004).
@@
In this paper, we choose $r_{g}=2GM_{\rm{bh}}/c^{2}$ as the unit of distance, $c$ as unit of velocity, $r_{g}c$ as unit of angular momentum, and $r_{g}/c$ as unit of time, unless specified otherwise. Here, $G$ is the gravitational constant, $M_{\rm{bh}}$ is the mass of the black hole and $c$ is speed of light in free space.
@@ 2.1 Simulation set-up
Our simulation set-up is very much similar to those which have been used over decades to study such shocked accretion flow (Chakrabarti and Molteni 1993; Ryu et al. 1995; Molteni et al. 1996a; Chakrabarti et al. 2004; Okuda et al. 2007; Giri et al. 2010; Okuda and Molteni 2012; Das et al. 2014; Lee et al. 2016; Kim et al. 2017, 2019; Garain and Kim 2023; Debnath et al. 2024). We perform our simulations on a vertically elongated RZR-Z domain $0:250]\times[{\color[rgb]{0,0,0}-2651,2651}]$. All the simulations are run using $480\times 2000$ ratioed grid points having a common ratio $1.0043$. Grids are constructed such that we resolve the region close to the origin with the highest resolution and grids become coarser as we move away. Thus, along the $R$ direction, finer grids are located close to $R=0$ and along the $Z$ direction, ratioing is done symmetrically about the equatorial plane. The black hole is placed at the origin, and we apply an absorbing boundary condition within a radius $r=1.8$. Mass of the black hole is chosen to be $_{\rm{bh}}=6.5\times 10^{9} M. M$ass injection rate $\dot{m}=0.01$ mass Eddington rate is assumed. We present the results of a total of five simulations here.
@@
Table 1: Simulation parameters where $l$, $v_{R0}$ and $a_{R0}$ represent the specific angular momentum of the accretion flow, radial velocity and sound speed of the injected matter respectively. See text for details.
Table 1 shows the simulation parameters presented in this paper. Column 1 shows the case IDs. Column 2 shows the specific angular momentum of the accretion flow $l$ at the outer boundary. Columns 3 and 4 show the radial velocity $v_{R0}$ and the sound speed $a_{R0}$ of the injected matter at the outer boundary. These are computed following the theoretical calculation provided in Chakrabarti (1989). The calculation requires two conserved parameters, specific energy $\epsilon=\frac{1}{2}(v_{R}^{2}+v_{\phi}^{2}+v_{Z}^{2})+\frac{a^{2}}{\gamma-1}+\Phi$ ($a=\frac{\gamma P}{\rho}$ being the local sound speed) and specific angular momentum $l$ of the flow, to figure out the radial variation of the flow profile. For all our runs, we choose $\epsilon=0.001$. This $\epsilon$, along with $l$ (Column 2), determines $v_{R0}$ and $a_{R0}$ at $R=R_{out}$. Our choice of ($l,\epsilon$) pairs is guided by the parameter space classification of this flow, Chakrabarti and Das (2001) which shows that for these sets, shocks can develop in the accretion solution. Additionally, we wish to investigate the effect of increased centrifugal barrier on the outflow and hence increase $l$ keeping $\epsilon$ constant. The last column represents the terminal velocities attained by the outflowing matter in the respective simulation. All simulations are run until a time stop of 160000. The solutions have achieved steady states by this time.
@@ 2.3 Bremsstrahlung emission modeling and self-Comptonization
Inside an accretion disk around a super-massive black hole, thermal bremsstrahlung (free–free emission) constitutes an important radiative mechanism contributing to X-ray and $\gamma$-ray emissions. Since we perform simulations of geometrically thick accretion disks ($h(R)\sim R$, where $h$ is the accretion disk half-height at radius $R$) with small $\dot{m}$, we have low optical depth values (optical depth of 0.15 cm-1 as shown in Section 3.3) Additionally, majority($\approx 99\%$) of the grid points in our simulation domain have temperature $2\times 10^{7}K$ hence we consider bremsstrahlung to be the dominant radiative mechanism and ignore line emissions. The spectral emissivity (erg/cc/sec/Hz) due to bremsstrahlung emission is given by (Rybicki and Lightman 1991):
@@
The seed photons for Comptonization are generated using the weighted photon distribution described above, followed by modeling the energy distribution through the Gamma function modeling (Pozdnyakov et al. 1983c; McGrath and Irving 1975) using a single random number.
@@ 3.1 Flow dynamics
Figure 1: shows the mass density distribution (in color) on log scale overlaid with velocity arrows at the final time for all cases A1-A5 (a:A1, b:A2, c:A3, d:A4, e:A5). Black color shows lower densities, and red color shows higher densities. Length of a velocity arrow is proportional to the magnitude $\sqrt{v_{R}^{2}+v_{Z}^{2}}$. Colors of arrows represent their magnitudes as shown in the legend for velocity vector. The maximum velocity is found extremely close to the black hole on the equatorial plane
@@
Figure 2: (a) shows a zoomed-in view of case A5 in the RZR-Z domain $0:150]\times[-160:160]$. We plot the number density (measured in number/$c$) colormap overlaid with velocity arrows. The iso-density contour (green line) corresponding to the value $8.85\times 10^{6}$, allows to identify the shock surface at $R\sim 100$. The region inside this contour is the high-density post-shock region, and the outer low-density region is the pre-shock region, marked accordingly. The post-shock region shows presence of turbulent vortices. We also mark the jet launching region close to the black hole. (b) This figure shows the temperature (in Kelvin) distribution in the same domain clearly separating the high-temperature post-shock region from the low-temperature pre-shock region for case A5.
@@
Zoomed-in observation of the equatorial region reveals the thick disk structure as well as the outflow launching mechanism. Fig  2 shows the central region $0:150]\times[-160:160]$ for the case A5. Fig  2(a) shows the number density distribution and (b) shows the temperature distribution. We notice that as the infalling matter moves towards the central region, density (and temperature as well) increases due to geometric compression initially and then due to a shock discontinuity. The discontinuity structure bends outward away from the equatorial plane (Molteni et al. 1994, 1996a) giving rise to a centrifugal pressure-supported boundary layer (CENBOL). Matter density increases past this discontinuity surface again due to geometric compression. An iso-density contour plot may allow one see the geometrically thick, toroidal disk structure in this region. The green iso-density contour corresponding to number density value $8.85\times 10^{6}$ in Fig  2(a) shows such a toroidal structure. Velocity vectors in the post-shock region show presence of turbulent vortices as well. We also notice from the velocity vectors that outflow is primarily launched from this post-shock region. This observation holds for all the cases. However, the centrifugal force varies from one case to another and that affects the shock strength and hence the compression. This causes the results to differ. We provide further analysis below to quantify these.
@@
Figure 3: represents the radial variations (along the equator) of logarithm of temperature($T$), number density ($n$), and velocity ($v_{R}$) at the final time. Different quantities are translated by a scalar along vertical axis to bring on same scale. We show temperature in Kelvin, number density in number/cc and velocity in cm/s units.
@@
The discontinuous jump and subsequent compression become prominent when we plot the radial variation of different fluid variables. Figure 3 shows the radial variation of number density, velocity, and temperature along the equatorial line for all the cases at the final time. Density and temperature are measured in particle number/cc and Kelvin units, respectively. For the low angular momentum cases ($\lambda=1.50$ and $1.60$, corresponding to runs A1 and A2 respectively), the flow shows a mild deceleration (velocity drop) at the centrifugal barrier around $R=28$ (A1) and $R=32$ (A2). A density compression (compression ratio $\rho_{\rm post-shock}/\rho_{\rm pre-shock}\sim 6.25$(A1) and $\sim 6.28$(A2)) and temperature enhancement is seen due to this. For slightly higher angular momentum $\lambda=1.70$ (run A3), the centrifugal barrier becomes strong enough to produce a well-defined shock (supersonic to subsonic transition) at $R=34.5$, resulting in a sharp, larger drop in velocity. This causes larger density compression and temperature rise. The density compression ratio is found to be $7.4$ in this case. We also observe the presence of an inner shock near $R=5$. For $\lambda=1.75$ (run A4), the shock feature is more pronounced (shock at $R=34.8$), and the density compression ratio is found to be $8.7$. The flow also exhibits evidence of a secondary, inner shock closer to the black hole ($R=13.5$) before passing through the inner sonic point. For the highest angular momentum $\lambda=1.80$ (run A5) case, the flow shows unsteady behavior, both radial and vertical oscillations of the post-shock matter are observed. Careful observation of successive snapshots show the formation and disappearance of turbulent vortices of different sizes in the post-shock matter. Since the centrifugal barrier is very strong in this case, the bounced back matter from this barrier interacts with the incoming matter, resulting in the initial formation of large vortices, above and below the equator. Subsequent velocity shear between the the incoming flow and the matter in these vortices develop shear instability. This instability possibly grows in time, resulting in the formation of turbulent flow. Thus, the post-shock region becomes highly turbulent with many intermediate discontinuities. The radial profiles of velocity, number density, and temperature at the final time for this case show such discontinuities. Such unsteady shock dynamics for high angular momentum flow have been reported in earlier numerical simulations (Ryu et al. 1997; Deb et al. 2016; Garain and Kim 2023).
@@ 3.2 Outflow
To investigate whether the outflowing matter can actually become unbounded, we plot the specific energy($\epsilon$) distribution inside the disk-outflow region in Fig. 5(a). These plots have been done using the final timestep data. For all the runs, infalling matter has $\epsilon=10^{-3}$. This value is mostly conserved within the accretion disk of thickness 280<Z<280-280<Z<280 at all radii except very close to the black hole. Inside the compressed region, energy is redistributed because of the turbulent interaction between the incoming and the centrifugal barrier reflected matter. Thus, we observe some matter even with a negative $\epsilon$. Such matter will be ultimately absorbed by the black hole. Importantly, we observe that the outflowing matter generally has higher values of the specific energy and is always positive, making sure that this matter will contribute to the unbound outflow.
Figure 5: (a) shows a colormap of specific energy distribution for all cases A1-A5 at the final time. The high-energy matter is shown by red color, which forms the outflow. (b) shows the force vectors overlaid on the density colormap for case A3, clearly showing unidirectional force vectors beyond 200-200 and $200$ in the Z-direction. Force is calculated in units of $\frac{c^{2}}{r_{g}}.$
@@
From Fig.  6, we observe that outflowing matter accelerates initially and attains a terminal velocity as it moves away. The initial acceleration to the outflowing matter is provided by the pressure gradient and centrifugal forces. We show the net force distribution (force vectors overlaid on number density distribution) around the central region for case A3 in Fig.  5(b). This plot reveals that the net force below $|Z|\lesssim 200$ is mostly inward. However, a very thin layer of matter close to the Z-axis experiences outward force. Because of the K-H instability in the region, it is very difficult to clearly identify unidirectional outward force. For $|Z|>200$, we start observing outward directed force vectors. This force imparts initial acceleration to the outflowing matter. However, further acceleration stops as matter moves away from the denser accretion disk. In the absence of any force away from the black hole, it attains a terminal velocity. The net force at the outflow base accelerates the matter beyond the escape velocity so that it could escape the gravitational pull of the central black hole. For low angular momentum runs (A1 and A2), the terminal velocity of the outflow remains small, saturating at $v_{z}\lesssim 0.09$. For these cases, density and pressure compressions at the discontinuity are lower, as can be seen from Fig.  3. Also, the centrifugal force is less due to lower values of $l$. For intermediate values ($\lambda=1.70$ run A3), the outflow velocity increases more efficiently, reaching $v_{z}\approx 0.12$. Here, the compressions and centrifugal force are higher than in the previous two cases. The maximum terminal velocity is obtained for $\lambda=1.75$ (run A4), where $v_{z}$ saturates near $0.14$. We observe the highest compressions in this case. For the highest angular momentum run ($\lambda=1.80$ run A5), the terminal velocity is significantly reduced ($\sim 0.1$). In this case, although the centrifugal force is highest, the compression is diluted due to several shocks and turbulence formations.

### 13 Apr 2026 v2

# 20XX Vol. X No. XX, 000–000

@@ Abstract
We study outflow from an unmagnetized, shocked accretion disk around a non-rotating super-massive black hole using multidimensional hydrodynamics simulation with radiative cooling. We aim to investigate whether such shocked accretion flow can launch sustained collimated bipolar outflow reaching out to thousands of gravitational radii even in the absence of magnetic field and if yes, what terminal velocity can they achieve? We present the results of a few simulations of geometrically thick accretion flow with increasing specific angular momentum on a vertically elongated cylindrical domain. We show that bipolar outflow from a region very close to the black hole is originating and propagating vertically out to our simulation domain boundary at around $2651$ Schwarzschild radius. The outflow attains a terminal velocity with a maximum value found to be $0.14c$ and the outflow rate depends on the angular momentum value of the accreting material. We also compute the self-Comptonized bremsstrahlung spectra for all the disk-jet runs.
@@ 1 Introduction
Many high-energy sources containing black hole show wind and/or jet (Mirabel and Rodríguez 1994; Stirling et al. 2001; Gallo 2010; Blandford et al. 2019). Whether the black hole is part of a binary system or located at the core of a galaxy, the qualitative structure of outflow is very similar when it is present. The outflow originates from very close to the central object and extends from hundreds to thousands of gravitational radii. It is well understood that, unlike a normal star, a black hole itself does not eject matter. Rather, the gravitationally captured matter, forming an accretion disk around the black hole, produces the outflow. Given that the gravitational attraction is very strong close to a black hole, it is thus very surprising how a fraction of the infalling matter gets rid of this strong attraction and becomes unbounded.
@@
Outflow formation from the magnetized plasma of an accretion disk is mainly attributed to two mechanisms: Blandford-Znayek (Blandford and Znajek 1977) and Blandford-Pyne (Blandford and Payne 1982). It is, of course not very surprising that the accreting material is magnetized. Magnetic field is very pervasive and wanders along with plasma. In a binary system, matter can drag in magnetic field from the companion star. In the case of accretion onto a galaxy center, the infalling matter can drag magnetic field from nearby stars or from the interstellar medium. As this magnetized matter moves closer to the black hole, the magnetic field structure, as well as the magnitude may change due to various physical processes such as geometric compression, dynamo effect, magnetic reconnection, etc. This magnetic field can then help in shaping the accretion disk structure and launching outflow from the accretion disk in the manner described in these references.
@@
However, the magnetic field may not be the only driving force behind launching outflow. In principle, if a sufficiently strong force develops in a way that cancels the gravitational pull, matter can be driven away from the accretion disk. Thus, if the outward centrifugal force due to rotating matter becomes sufficiently strong, the effective potential barrier (Misner et al. 1973; Shapiro and Teukolsky 1983) may be so high that the matter may not be able to accrete onto the central black hole and be driven away. Also, a sufficiently strong pressure gradient force acting opposite to the gravitational attraction can drive material away. The pressure can be due to radiation or thermal. If the disk material and the radiation are interacting effectively, radiation pressure can act in that case (Icke 1989; Fukue 1996; Chattopadhyay et al. 2004; Vyas et al. 2015). On the other hand, if the cooling time scale is lower than the infall time scale, matter can become sufficiently hot and the required thermal pressure gradient force can develop inside the accretion flow (e.g., Blandford and Begelman 1999; Soker 2023). If an accretion disk, having satisfied such conditions, additionally carries magnetic field, all the effects may add up to make the outflow launching favorable (e.g., Fukue 2004).
@@
In this paper, we choose $r_{g}=2GM_{\rm{bh}}/c^{2}$ as the unit of distance, $c$ as unit of velocity, $r_{g}c$ as unit of angular momentum, and $r_{g}/c$ as unit of time, unless specified otherwise. Here, $G$ is the gravitational constant, $M_{\rm{bh}}$ is the mass of the black hole, and $c$ is the speed of light in free space.
@@ 2.1 Simulation set-up
Our simulation set-up is very much similar to those which have been used over decades to study such shocked accretion flow (Chakrabarti and Molteni 1993; Ryu et al. 1995; Molteni et al. 1996a; Chakrabarti et al. 2004; Okuda et al. 2007; Giri et al. 2010; Okuda and Molteni 2012; Das et al. 2014; Lee et al. 2016; Kim et al. 2017, 2019; Garain and Kim 2023; Debnath et al. 2024). We perform our simulations on a vertically elongated RZR-Z domain $0:250]\times[-2651:2651]$. All the simulations are run using $480\times 2000$ ratioed grid points having a common ratio $1.0043$. Grids are constructed such that we resolve the region close to the origin with the highest resolution and grids become coarser as we move away. Thus, along the $R$ direction, finer grids are located close to $R=0$ and along the $Z$ direction, ratioing is done symmetrically about the equatorial plane. The black hole is placed at the origin, and we apply an absorbing boundary condition within a radius $r=1.8$. Mass of the black hole is chosen to be $_{\rm{bh}}=6.5\times 10^{9} M. M$ass injection rate $\dot{m}=0.01$ mass Eddington rate is assumed. We present the results of a total of five simulations here.
@@
Table 1: Simulation parameters where $l$, $v_{R0}$ and $a_{R0}$ represent the specific angular momentum of the accretion flow, radial velocity and sound speed of the injected matter respectively. Last column represents the terminal velocity of the outflow. See text for details.
Table 1 shows the simulation parameters presented in this paper. Column 1 shows the case IDs. Column 2 shows the specific angular momentum of the accretion flow $l$ at the outer boundary. Columns 3 and 4 show the radial velocity $v_{R0}$ and the sound speed $a_{R0}$ of the injected matter at the outer boundary. These are computed following the theoretical calculation provided in Chakrabarti (1989). The calculation requires two conserved parameters, specific energy $\epsilon=\frac{1}{2}(v_{R}^{2}+v_{\phi}^{2}+v_{Z}^{2})+\frac{a^{2}}{\gamma-1}+\Phi$ ($a=\frac{\gamma P}{\rho}$ being the local sound speed) and specific angular momentum $l$ of the flow, to figure out the radial variation of the flow profile. For all our runs, we choose $\epsilon=0.001$. This $\epsilon$, along with $l$ (Column 2), determines $v_{R0}$ and $a_{R0}$ at $R=R_{out}$. Our choice of ($l,\epsilon$) is guided by the parameter space classification of this flow, Chakrabarti and Das (2001) which shows that for these sets, shocks can develop in the accretion solution. Additionally, we wish to investigate the effect of increased centrifugal barrier on the outflow and hence increase $l$ keeping $\epsilon$ constant. The last column represents the terminal velocities attained by the outflowing matter in the respective simulation. All simulations are run until a time stop of 160000. The solutions have achieved steady states by this time.
@@ 2.3 Bremsstrahlung emission modeling and self-Comptonization
Inside an accretion disk around a super-massive black hole, thermal bremsstrahlung (free–free emission) constitutes an important radiative mechanism contributing to X-ray and $\gamma$-ray emissions. Since we perform simulations of geometrically thick accretion disks ($h(R)\sim R$, where $h$ is the accretion disk half-height at radius $R$) with small $\dot{m}$, we have low optical depth values (optical depth of 0.15 cm-1 as shown in Section 3.3) Additionally, majority($\approx 99\%$) of the grid points in our simulation domain have temperature $2\times 10^{7}K$, hence we consider bremsstrahlung to be the dominant radiative mechanism and ignore line emissions. The spectral emissivity (erg/cc/sec/Hz) due to bremsstrahlung emission is given by (Rybicki and Lightman 1991):
@@
The seed photons for Comptonization are generated using the weighted photon distribution described above, followed by modeling the energy distribution through Gamma function modeling (Pozdnyakov et al. 1983c; McGrath and Irving 1975) using a single random number.
@@ 3.1 Flow dynamics
Figure 1: shows the mass density ($\rho$) distribution (in color) on log scale overlaid with velocity arrows at the final time for all cases A1-A5 (a:A1, b:A2, c:A3, d:A4, e:A5). Black color shows lower densities, and red color shows higher densities. Length of a velocity arrow is proportional to the magnitude $\sqrt{v_{R}^{2}+v_{Z}^{2}}$. Colors of arrows represent their magnitudes as shown in the legend for velocity vector. The maximum velocity is found extremely close to the black hole on the equatorial plane
@@
Figure 2: (a) shows a zoomed-in view of case A5 in the RZR-Z domain $0:150]\times[-160:160]$. We plot the number density (measured in particle number/$c$) colormap overlaid with velocity arrows. The iso-density contour (green line) corresponding to the value $8.85\times 10^{6}$, allows to identify the shock surface at $R\sim 100$. The region inside this contour is the high-density post-shock region, and the outer low-density region is the pre-shock region, marked accordingly. The post-shock region shows the presence of turbulent vortices. We also mark the jet launching region close to the black hole. (b) This figure shows the temperature (in Kelvin) distribution in the same domain clearly separating the high-temperature post-shock region from the low-temperature pre-shock region for case A5.
@@
Zoomed-in observation of the equatorial region reveals the thick disk structure as well as the outflow launching mechanism. Fig  2 shows the central region $0:150]\times[-160:160]$ for the case A5. Fig  2(a) shows the number density distribution and (b) shows the temperature distribution. We notice that as the infalling matter moves towards the central region, density (and temperature as well) increases due to geometric compression initially and then due to a shock discontinuity. The discontinuity structure bends outward away from the equatorial plane (Molteni et al. 1994, 1996a) giving rise to a centrifugal pressure-supported boundary layer (CENBOL). Matter density increases past this discontinuity surface again due to geometric compression. An iso-density contour plot may allow one to see the geometrically thick, toroidal disk structure in this region. The green iso-density contour corresponding to number density value $8.85\times 10^{6}$ in Fig  2(a) shows such a toroidal structure. Velocity vectors in the post-shock region show presence of turbulent vortices as well. We also notice from the velocity vectors that outflow is primarily launched from this post-shock region. This observation holds for all the cases. However, the centrifugal force varies from one case to another and that affects the shock strength and, hence, the compression. This causes the results to differ. We provide further analysis below to quantify these.
@@
Figure 3: represents the radial variations (along the equator) of logarithm of temperature($T$), number density ($n$), and velocity ($v_{R}$) at the final time. Different quantities are translated by a scalar along the vertical axis to bring them to the same scale. We show temperature in Kelvin, number density in particle number/cc and velocity in cm/s units.
@@
The discontinuous jump and subsequent compression become prominent when we plot the radial variation of different fluid variables. Figure 3 shows the radial variation of number density, velocity, and temperature along the equatorial line for all the cases at the final time. Density and temperature are measured in particle number/cc and Kelvin units, respectively. For the low angular momentum cases ($\lambda=1.50$ and $1.60$, corresponding to runs A1 and A2 respectively), the flow shows a mild deceleration (velocity drop) at the centrifugal barrier around $R=28$ (A1) and $R=32$ (A2). A density compression (compression ratio $\rho_{\rm post-shock}/\rho_{\rm pre-shock}\sim 6.25$(A1) and $\sim 6.28$(A2)) and temperature enhancement is seen due to this. For slightly higher angular momentum $\lambda=1.70$ (run A3), the centrifugal barrier becomes strong enough to produce a well-defined shock (supersonic to subsonic transition) at $R=34.5$, resulting in a sharp, larger drop in velocity. This causes larger density compression and temperature rise. The density compression ratio is found to be $7.4$ in this case. We also observe the presence of an inner shock near $R=5$. For $\lambda=1.75$ (run A4), the shock feature is more pronounced (shock at $R=34.8$), and the density compression ratio is found to be $8.7$. The flow also exhibits evidence of a secondary, inner shock closer to the black hole ($R=13.5$) before passing through the inner sonic point. For the highest angular momentum $\lambda=1.80$ (run A5) case, the flow shows unsteady behavior, both radial and vertical oscillations of the post-shock matter are observed. Careful observation of successive snapshots show the formation and disappearance of turbulent vortices of different sizes in the post-shock matter. Since the centrifugal barrier is very strong in this case, the bounced-back matter from this barrier interacts with the incoming matter, resulting in the initial formation of large vortices, above and below the equator. Subsequent velocity shear between the the incoming flow and the matter in these vortices develop shear instability. This instability possibly grows in time, resulting in the formation of turbulent flow. Thus, the post-shock region becomes highly turbulent with many intermediate discontinuities. The radial profiles of velocity, number density, and temperature at the final time for this case show such discontinuities. Such unsteady shock dynamics for high angular momentum flow have been reported in earlier numerical simulations (Ryu et al. 1997; Deb et al. 2016; Garain and Kim 2023).
@@ 3.2 Outflow
To investigate whether the outflowing matter can actually become unbounded, we plot the specific energy($\epsilon$) distribution inside the disk-outflow region in Fig. 5(a). These plots have been done using the final timestep data. For all the runs, infalling matter has $\epsilon=10^{-3}$. This value is mostly conserved within the accretion disk of thickness 280<Z<280-280<Z<280 at all radii except very close to the black hole. Inside the compressed region, energy is redistributed because of the turbulent interaction between the incoming and the centrifugal barrier reflected matter. Thus, we observe some matter even with a negative $\epsilon$. Such matter will be ultimately absorbed by the black hole. Importantly, we observe that the outflowing matter generally has higher values of the $\epsilon$ and is always positive, making sure that this matter will contribute to the unbound outflow.
[…]
@@
[…]